How do I explain in one paragraph how to simplify 16 1/2?

Found 3 solutions by OkayImaknowitall23, Alan3354, lynnlo:
Answer by OkayImaknowitall23(5)

In one paragraph you could:
The equation 16 1/2 can simplify by first turning the mixed number into a improper fraction which you can do by multiplying the whole number (16) into the denominator (2) which equals to 32 the add it to the numerator 1. It will give a fraction stating : 32/2 simplified by diving the denominator and the numerator by 2 and you get 16/1

Answer by Alan3354(69443)

How do I explain in one paragraph how to simplify 16 1/2?
--------------
Simplified can be a matter of opinion.
16 1/2 = 16.5
16 1/2 = 33/2
Which is simpler?

Answer by lynnlo(4176)

HOW TO SIMPLIFY 16 1/2
FIRST OF ALL,YOU KNOW THIS IS A COMPOUND FRACTION.
IN THIS PROBLEM YOU WILL NEED TO MULTIPLY AND ADD.
THINK OF YOUR PROBLEM AS JUST ADDING TWO NUMBERS TOGETHER,WHEN YOU SAY 16 AND A HALF,YOU WILL USE THE CONJUNCTION,FIRST YOU WILL HAVE TO HAVE THE SAME DENOMINATORS BEFORE YOU CAN ADD THEM TOGETHER.YOU WILL HAVE TO FIND THE LEAST COMMON DENOMINATOR.
16(1)=16*(2/2)=(16*2)/2=32/2
THEN ADD YOUR FRACTIONS
16+1/2=32/2+1/2=33/2
SIMPLIFIED FRACTION IS 33/2